OpenStudy (anonymous):

(3/5) to the 3rd power times (1/3) =

OpenStudy (kc_kennylau):

Write (3/5) to the 3rd power as (3/5)(3/5)(3/5), and then just multiply all the denominators together and all the numerators together.

OpenStudy (anonymous):

runner5000: http://openstudy.com/updates/53506acbe4b0285dde94dd30 will give medals and will be fanned please help me for god sakes going to every question for someone to help me out please help

OpenStudy (anonymous):

so it would be 27/36?

OpenStudy (kc_kennylau):

Well, I shall do it step by step for you, and you'll tell me which step you don't understand if there's any

OpenStudy (anonymous):

c1282 help me out please

OpenStudy (anonymous):

no i get it you mutiple it and it should be 27/12 times 1/3 which equals 27/36

OpenStudy (kc_kennylau):

Firstly rewrite the question using notation: \(\left(\dfrac35\right)^3\left(\dfrac13\right)\) Then expand the index: \(\left(\dfrac35\right)\left(\dfrac35\right)\left(\dfrac35\right)\left(\dfrac13\right)\) Then multiply the denominators together and the numerators together: \(\dfrac{3\times3\times3\times1}{5\times5\times5\times3}\) Then evaluate the value:\(\dfrac{27}{375}\)

OpenStudy (kc_kennylau):

Can you tell me how you got 27/12 times 1/3? I want to help you by looking at what you did wrong.

OpenStudy (anonymous):

i only multiplied the first fraction then multipled it by the second one

OpenStudy (kc_kennylau):

Because arithmetic is the foundation of mathematics

OpenStudy (kc_kennylau):

Wait... oh so that's a typo, 27/125 instead of 27/12

OpenStudy (kc_kennylau):

Is that what you mean?

OpenStudy (anonymous):

i did 3 x 3 x3 then did 5 x 5 x 5 that equals 27/12

OpenStudy (anonymous):

then i took that and multipled it by 1/3

OpenStudy (kc_kennylau):

You did it by hand or by a calculator?

OpenStudy (anonymous):

calulator

OpenStudy (kc_kennylau):

Because 5 x 5 x 5 equals 125 and I'm wondering where the missing digit would have been

OpenStudy (anonymous):

i got it now after ebing simplified it should 9/125

OpenStudy (anonymous):

being

OpenStudy (kc_kennylau):

Anyway, I don't think one should depend on a calculator. It's not hard to calculate 5 x 5 x 5 by heart I think? 5 x 5 x 5 = 25 x 5 = 20x5 + 5x5 = 100 + 25 = 125

OpenStudy (kc_kennylau):

Yep that's correct

OpenStudy (anonymous):

thank you

OpenStudy (kc_kennylau):

no problem

OpenStudy (anonymous):

runner5000: http://openstudy.com/updates/53506acbe4b0285dde94dd30 will give medals and will be fanned please help me out please will give medals